# The Cosmological CPT Theorem

### v2

@@ 3.3 Inflationary Observables
In the cosmological context the purpose of studying de Sitter spacetime is to understand the behaviour of quantum fluctuations at the end of inflation. These perturbations then grow over the history of the universe to produce all the structure that we see. Due to the randomness inherent to the quantum nature of these fluctuations and the single universe that we have to observe we study them through their correlation functions. In this section we focus on cosmological correlators calculated using the in-in formalism, since constraints for them only require the spacetime symmetries considered in Section 3.2. In this paper, we will focus only on observables involving fields with integer spin representations and will not consider spinor fields in detail. However, a similar formalism should provide a much easier approach to deriving symmetry and Unitarity constraints for observables involving half-integer spin fields in cosmology (see e.g. [106, 107, 108, 109, 110, 111, 112, 113, 114, 86, 115, 116, 117, 118, 119] for discussions of spinors in de Sitter space and cosmology), which has yet to be done in the cosmology literature.
@@
$\langle\mathcal{O}(\eta)\rangle=\langle\Psi(\eta)\rvert\mathcal{O}(\eta)\lvert\Psi(\eta)\rangle\,.$ (3.53)
We do not (a priori) know the state $\lvert\Psi(\eta)\rangle$. However, we assume that the universe was in a vacuum of the full theory, $\lvert\Psi\rangle=\lvert\Omega\rangle$. Just like in flat space it will prove very helpful to move to the interaction picture where operators evolve according to the free Hamiltonian whilst the evolution of states is dictated by the interacting Hamiltonian. Then these correlators are
@@
However, $\lvert\Omega(\eta_{\text{i}})\rangle$ is the in vacuum of the full theory (despite being expressed in the interaction picture). This is relevant as the operators, $\mathcal{O}$, will be constructed from the interaction picture creation and annihilation operators, which act straightforwardly on the free (Bunch-Davies) vacuum, |0$\lvert$ 0$\rangle$| 0 ⟩. By taking $\eta_{i}$ to the infinite past we can turn off the interactions adiabatically so that the free and interacting vacua coincide. We will discuss the procedure by which this is done more extensively in Section 7.1.
@@ 4.1 Distinction from Spacetime Symmetries
    A spacetime symmetry, which acts on the entire manifold $\cal M$ by moving points around by means of some map $m:x\to y$. (In addition to some local transformation of the fields $\phi(x)$, and possibly a complex conjugation of the amplitude.)
@@
There can also exist local Lagrangian symmetries which do not arise from any spacetime symmetry of $\cal M$. A classic example of this, is the standard assumption that a field theory has local Lorentz Invariance at each point $p$, even on a spacetime manifold which has no Killing symmetry. (This example also shows that the definition of “local” has nothing to do with gauge invariance, as we typically assume local Lorentz Invariance even when doing field theory on a fixed, non-gravitational background.) Another simple example would be a time-varying Hamiltonian $H(t)$, in a case where at each moment of time $t$ the Hamiltonian is time-reversal invariant, but there exists no time $t_{0}$ for which $H(t_{0}+t)=H(t_{0}-t)$.
Let us now consider the implications for the discrete symmetries in our $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group. In the case of Reflection Reality ($\bf{RR}$), the fixed points include the entire real part of the Lorentzian manifold $\cal M$. Hence, if we are working in Lorentzian signature, we are automatically entitled to extend $\bf{RR}$ to a symmetry of the local Lagrangian at each point $p$.
@@ 4.2 Flat Space Wavefunction
Figure 1: The vacuum wavefunction $\Psi[\tau=0]$ is given by a path integral over one half of the Euclidean $\tau$ plane, with the usual conventions integrating over $-\infty<\tau\leq 0$.
@@
[…]
@@
where we have used the fact that (as discussed in Section 3.1) for equal-time quantities, such as the de Sitter wavefunction, $\bf{CRT}$ simplifies to a complex conjugation, a flip in the global time coordinate and a rotation on the $d-1$-sphere. Note that at $T=0$, $\bf{CRT}$ acts to constrain the wavefunction to be its own conjugate after a spatial reflection.
@@ 4.3 Poincaré Wavefunction of the Universe
Poincaré $\bf{CRT}$ + $SO^{+}(d+1,1)$implies $\mathbf{LL}$ $\bf{CRT}$.
@@ 5.1 Perturbative Equivalence of Reflection Reality and Unitarity
Suppose we start with a unitary QFT, and consider deforming it by some small perturbation $\delta L$ to the Lagrangian density $L$ [^35] Even if a QFT does not come from a Lagrangian, perturbations to the theory still do. Let us assume that this small perturbation does not change the number of configuration degrees of freedom in the theory. [^36] If the perturbation involves higher derivative couplings, it may be necessary to assume it comes with powers of the UV cutoff, to prevent additional degrees of freedom from appearing. Throughout this section, we are working in the regime of naive QFT perturbation theory, where we assume the Hilbert space is independent of the choice of coupling constants. This is not really true, but we expect that the conclusions of this section would not be affected by a more sophisticated analysis. It should then be possible to view it as a change to the Hamiltonian $\delta H$. If we consider all possible terms in the Hamiltonian (satisfying some set of symmetries) up to some degree of irrelevance, this should give us a finite-dimensional space of perturbations to the couplings, so $H\in\mathbb{C}^{N}$ for some finite dimension $N$. Now the subspace of perturbations that preserve Unitarity should be those satisfying:
@@
For a bosonic field theory in Lorentzian signature, such perturbations are in one-to-one correspondence with real contributions to $\delta L$ [^37] There is, indeed, one class of imaginary contributions to $\delta L$ that are compatible with Unitarity, namely imaginary total derivatives. These correspond to imaginary canonical transformations of the state space, which can be viewed as rescaling the definition of the inner product. But because the action appears inside of an exponential, such rescalings are multiplicative and thus cannot spoil Unitarity. If however, we consider perturbations that involve derivative couplings, in Lorentzian signature there can also be (typically divergent) measure factors required to restore Unitarity, that are tantamount to fixing certain imaginary terms in the action. However, upon passing to Euclidean signature, such terms are fixed by $\bf{RR}$ alone, without using any additional feature of Unitarity. This is one way in which Euclidean field theory is nicer than Lorentzian field theory.
@@ 5.2 How to get Theories with Negative Norm States
In $D$ spacetime dimensions, the number of propagating degrees of freedom is $D-1$ when $m^{2}\neq 0$, but only $D-2$ when $m^{2}=0$. The latter is, of course, due to the usual gauge symmetry of a massless photon field. Since the gauge theory has a bunch of null sates, which have 0 inner product with any other vector, it can be a boundary between a unitary theory, and a theory containing negative norm states. And in fact, in the tachyonic case where $m^{2}<0$, there is a propagating temporal mode which has negative norm. If then, you have a unitary theory with a gauge symmetry, and you want to deform it to another unitary theory, the safest thing to do is to restrict attention to only gauge-invariant perturbations to the Lagrangian.
@@ 6 Analytic Continuation in Cosmology
Similar to how the physical domain of Mandelstam variables in amplitudes is defined for $s,t\geq 0$ and crossing symmetry involves analytically continuing to $s,t\in\mathbb{C}$ (see [123] for recent progress regarding crossing in flat space amplitudes), we will now discuss how our discrete transformations in Section 3.2 will involve analytically continuing the conformal time outside of the physical domain of $\eta\leq 0$ and in Fourier space will also involve analytically continuing the magnitude of the momenta $k=\sqrt{{\textbf{k}}\cdot{\textbf{k}}}\geq 0$.
@@ 6.1 Non-perturbative Analytic Continuation
  • $A=-i$, $B=+1$: Past dS-Poincaré $P^{-}$, with $\zeta=-t$;
@@
  • $A=B=-i$: Funky Euclidean hyperbolic space $H^{-}$ with “all past” minus signature, with $\zeta=\tau$.
@@
At $q=\tfrac{1}{2}$, this contour passes through funky hyperbolic space $H^{+}$ where the QFT need not converge. Worse still, at $q=1$, it ends up at $A=+i$, $B=-1$ which is not the same as $P^{-}$!
@@
${\cal A}^{-}(\eta;y_{1},\dots,y_{d};\Omega)={\cal A}^{+}(e^{-i\pi}\eta;y_{1},\dots,y_{d};e^{-i\pi}\Omega)\,,$ (6.5)
where in this expression, the use of $e^{i\pi}$ rather than 1-1- 1 is a mnemonic indicating in which direction one should rotate in the complex plane to reach 1-1- 1 (all such rotations are to be done simultaneously, in this case holding the ratio of $\eta$ and $\Omega$ fixed). But $P^{+}$ and $P^{-}$ are also related by $\bf{CRT}$ symmetry, which tells us that $P^{+}$ and $P^{-}$ are also related by sending $y_{i}\to-y_{i}$ (which is explicitly the reflection $\mathbf{R}$) and complex conjugating the amplitude. Hence, invariance under $\bf{CRT}$ also requires $P^{+}$ and $P^{-}$ to also be related by the transformation:
${\cal A}^{-}(\eta;y_{1},\dots,y_{d};\Omega)={\cal A}^{+}(\eta;-y_{1},\dots,-y_{d};\Omega)^{*}\,.$ (6.6)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;y_{1},\dots,y_{d};\Omega)=\psi_{n}(e^{-i\pi}\eta;-y_{1},\dots,-y_{d};e^{-i\pi}\Omega)\,.$ (6.7)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;k_{1},\dots,k_{d};\Omega)=\psi_{n}(e^{-i\pi}\eta;k_{1},\dots,k_{d};e^{-i\pi}\Omega)\,,$ (6.8)
@@
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;y_{1},\dots,y_{d};\Omega)=\psi_{n}(e^{\pm i\pi}\eta;e^{\pm i\pi}y_{1},\dots,e^{\pm i\pi}y_{d};\Omega)\,,$ (6.9)
@@
$\mathbf{RR}:\psi_{n}^{*}(\eta;y_{1},\dots,y_{d};\Omega)=\psi_{n}(\eta;-e^{i\pi}y_{1},\dots,-e^{i\pi}y_{d};e^{-i\pi}\Omega)\,,$ (6.10)
where, similarly to previous expressions, multiplication by $-e^{i\pi}$ is shorthand for an analytic continuation followed by a reflection, and cannot validly be replaced with +1+1+ 1!
@@
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;k_{1},\dots,k_{d};\Omega)=(-1)^{\pm d}\psi_{n}(e^{\pm i\pi}\eta;e^{\mp i\pi}k_{1},\dots,e^{\mp i\pi}k_{d};\Omega)\,,$ (6.11)
@@
$\mathbf{RR}:\psi_{n}^{*}(\eta;k_{1},\dots,k_{d};\Omega)=(-1)^{d}\psi_{n}(\eta;e^{-i\pi}k_{1},\dots,e^{-i\pi}k_{d};e^{i\pi}\Omega)\,.$ (6.12)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;y_{1},\dots,y_{d};\Omega)=$ $\psi_{n}(e^{-i\pi}\eta;-y_{1},\dots,-y_{d};e^{-i\pi}\Omega)\,;$ (6.13)
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;y_{1},\dots,y_{d};\Omega)=$ $\psi_{n}(e^{\pm i\pi}\eta;e^{\pm i\pi}y_{1},\dots,e^{\pm i\pi}y_{d};\Omega)\,;$ (6.14)
$\mathbf{RR}:\psi_{n}^{*}(\eta;y_{1},\dots,y_{d};\Omega)=$ $\psi_{n}(\eta;-e^{i\pi}y_{1},\dots,-e^{i\pi}y_{d};e^{-i\pi}\Omega)\,.$ (6.15)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;k_{1},\dots,k_{d};\Omega)=$ $\psi_{n}(e^{-i\pi}\eta;k_{1},\dots,k_{d};e^{-i\pi}\Omega)\,;$ (6.16)
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;k_{1},\dots,k_{d};\Omega)=$ $e^{\pm i\pi d}\,\psi_{n}(e^{\pm i\pi}\eta;e^{\mp i\pi}k_{1},\dots,e^{\mp i\pi}k_{d};\Omega)\,;$ (6.17)
$\mathbf{RR}:\psi_{n}^{*}(\eta;k_{1},\dots,k_{d};\Omega)=$ $e^{i\pi d}\,\psi_{n}(\eta;e^{-i\pi}k_{1},\dots,e^{-i\pi}k_{d};e^{-i\pi}\Omega)\,.$ (6.18)

### v3

@@ 3.3 Inflationary Observables
In the cosmological context the purpose of studying de Sitter spacetime is to understand the behaviour of quantum fluctuations at the end of inflation. These perturbations then grow over the history of the universe to produce all the structure that we see. Due to the randomness inherent to the quantum nature of these fluctuations and the single universe that we have to observe we study them through their correlation functions. In this section we focus on cosmological correlators calculated using the in-in formalism, since constraints for them only require the spacetime symmetries considered in Section 3.2. In this paper, we will focus only on observables involving fields with integer spin representations and will not consider spinor fields in detail. However, a similar formalism should provide a much easier approach to deriving symmetry and Unitarity constraints for observables involving half-integer spin fields in cosmology (see e.g. [111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 86, 121, 122, 123, 124, 125] for discussions of spinors in de Sitter space and cosmology), which has yet to be done in the cosmology literature.
@@
$\langle\mathcal{O}(\eta)\rangle=\langle\Psi\rvert\mathcal{O}(\eta)\lvert\Psi\rangle\,.$ (3.53)
We do not (a priori) know the state $\lvert\Psi\rangle$. However, we assume that the universe was in a vacuum of the full theory, $\lvert\Psi\rangle=\lvert\Omega\rangle$. Just like in flat space it will prove very helpful to move to the interaction picture where operators evolve according to the free Hamiltonian whilst the evolution of states is dictated by the interacting Hamiltonian. Then these correlators are
@@
However, $\lvert\Omega(\eta_{\text{i}})\rangle$ is the in vacuum of the full theory (despite being expressed in the interaction picture). This is relevant as the operators, $\mathcal{O}$, will be constructed from the interaction picture creation and annihilation operators, which act straightforwardly on the free (Bunch-Davies) vacuum, $\lvert 0\rangle$. By taking $\eta_{i}$ to the infinite past we can turn off the interactions adiabatically so that the free and interacting vacua coincide. We will discuss the procedure by which this is done more extensively in Section 7.1.
@@ 4.1 Distinction from Spacetime Symmetries
[…]
@@
[…]
@@ 4.2 Flat Space Wavefunction
Figure 1: The vacuum wavefunction $\Psi[\tau=0]$ is given by a path integral over one half of the Euclidean $\tau$ plane, with the usual conventions integrating over $\infty<\tau\leq 0$.
@@
The story above also applies, without any essential change, to global de Sitter where $\bf{CRT}$ allows us to write a concise non-perturbative constraint for the full de Sitter wavefunction at equal-$T$:
@@
where we have used the fact that (as discussed in Section 3.1) for equal-time quantities, such as the de Sitter wavefunction, $\bf{CRT}$ simplifies to a complex conjugation, a flip in the global time coordinate and a rotation on the d1d-1-sphere. Note that at $T=0$, $\bf{CRT}$ acts to constrain the wavefunction to be its own conjugate after a spatial reflection.
@@ 4.3 Poincaré Wavefunction of the Universe
Poincaré $\bf{CRT}$ + $SO^{+}(d+1,1)$ $\implies$ $\mathbf{LL}$ $\bf{CRT}$.
@@ 5.1 Perturbative Equivalence of Reflection Reality and Unitarity
Suppose we start with a unitary QFT, and consider deforming it by some small perturbation $\delta L$ to the Lagrangian density LL. [^35] Even if a QFT does not come from a Lagrangian, perturbations to the theory still do. Let us assume that this small perturbation does not change the number of configuration degrees of freedom in the theory. [^36] If the perturbation involves higher derivative couplings, it may be necessary to assume it comes with powers of the UV cutoff, to prevent additional degrees of freedom from appearing. Throughout this section, we are working in the regime of naive QFT perturbation theory, where we assume the Hilbert space is independent of the choice of coupling constants. This is not really true, but we expect that the conclusions of this section would not be affected by a more sophisticated analysis. It should then be possible to view it as a change to the Hamiltonian $\delta H$. If we consider all possible terms in the Hamiltonian (satisfying some set of symmetries) up to some degree of irrelevance, this should give us a finite-dimensional space of perturbations to the couplings, so $H\in\mathbb{C}^{N}$ for some finite dimension $N$. Now the subspace of perturbations that preserve Unitarity should be those satisfying:
@@
For a bosonic field theory in Lorentzian signature, such perturbations are in one-to-one correspondence with real contributions to $\delta L$. [^37] There is, indeed, one class of imaginary contributions to $\delta L$ that are compatible with Unitarity, namely imaginary total derivatives. These correspond to imaginary canonical transformations of the state space, which can be viewed as rescaling the definition of the inner product. But because the action appears inside of an exponential, such rescalings are multiplicative and thus cannot spoil Unitarity. If however, we consider perturbations that involve derivative couplings, in Lorentzian signature there can also be (typically divergent) measure factors required to restore Unitarity, that are tantamount to fixing certain imaginary terms in the action. However, upon passing to Euclidean signature, such terms are fixed by $\bf{RR}$ alone, without using any additional feature of Unitarity. This is one way in which Euclidean field theory is nicer than Lorentzian field theory.
@@ 5.2 How to get Theories with Negative Norm States
In $D$ spacetime dimensions, the number of propagating degrees of freedom is D1D-1 when $m^{2}\neq 0$, but only D2D-2 when $m^{2}=0$. The latter is, of course, due to the usual gauge symmetry of a massless photon field. Since the gauge theory has a bunch of null sates, which have 0 inner product with any other vector, it can be a boundary between a unitary theory, and a theory containing negative norm states. And in fact, in the tachyonic case where $m^{2}<0$, there is a propagating temporal mode which has negative norm. If then, you have a unitary theory with a gauge symmetry, and you want to deform it to another unitary theory, the safest thing to do is to restrict attention to only gauge-invariant perturbations to the Lagrangian.
@@ 6 Analytic Continuation in Cosmology
Similar to how the physical domain of Mandelstam variables in amplitudes is defined for $s,t\geq 0$ and crossing symmetry involves analytically continuing to $s,t\in\mathbb{C}$ (see [129] for recent progress regarding crossing in flat space amplitudes), we will now discuss how our discrete transformations in Section 3.2 will involve analytically continuing the conformal time outside of the physical domain of $\eta\leq 0$ and in Fourier space will also involve analytically continuing the magnitude of the momenta $k=\sqrt{{\textbf{k}}\cdot{\textbf{k}}}\geq 0$. We note that some similar, yet distinct, analytic continuations in the in-in formalism were performed in e.g. [130, 131, 132, 133, 134, 135, 136, 137, 138, 139].
@@ 6.1 Non-perturbative Analytic Continuation
  • A=iA=-i, $B=+1$: Past dS-Poincaré $P^{-}$, with $\zeta=-t$;
@@
  • A=B=iA=B=-i: Funky Euclidean hyperbolic space $H^{-}$ with “all past” minus signature, with $\zeta=\tau$.
@@
At $q=\tfrac{1}{2}$, this contour passes through funky hyperbolic space $H^{+}$ where the QFT need not converge. Worse still, at $q=1$, it ends up at $A=+i$, B=1B=-1 which is not the same as $P^{-}$!
@@
${\cal A}^{-}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)={\cal A}^{+}(e^{-i\pi}\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};e^{-i\pi}\Omega)\,$, (6.5)
where in this expression, the use of $e^{i\pi}$ rather than 1-1 is a mnemonic indicating in which direction one should rotate in the complex plane to reach 1-1 (all such rotations are to be done simultaneously, in this case holding the ratio of $\eta$ and $\Omega$ fixed). But $P^{+}$ and $P^{-}$ are also related by $\bf{CRT}$ symmetry, which tells us that $P^{+}$ and $P^{-}$ are also related by sending $y_{i}\to-y_{i}$ (which is explicitly the reflection $\mathbf{R}$) and complex conjugating the amplitude. Hence, invariance under $\bf{CRT}$ also requires $P^{+}$ and $P^{-}$ to also be related by the transformation:
${\cal A}^{-}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)={\cal A}^{+}(\eta;-\textbf{y}_{1},\dots,-\textbf{y}_{n};\Omega)^{*}\,.$ (6.6)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=\psi_{n}(e^{-i\pi}\eta;-\textbf{y}_{1},\dots,-\textbf{y}_{n};e^{-i\pi}\Omega)\,.$ (6.7)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=\psi_{n}(e^{-i\pi}\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};e^{-i\pi}\Omega)\,$, (6.8)
@@
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=\psi_{n}(e^{\pm i\pi}\eta;e^{\pm i\pi}\textbf{y}_{1},\dots,e^{\pm i\pi}\textbf{y}_{n};\Omega)\,$, (6.9)
@@
$\mathbf{RR}:\psi_{n}^{*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=\psi_{n}(\eta;-e^{i\pi}\textbf{y}_{1},\dots,e^{i\pi}\textbf{y}_{n};e^{-i\pi}\Omega)\,$, (6.10)
where, similarly to previous expressions, multiplication by $e^{i\pi}$ is shorthand for an analytic continuation followed by a reflection, and cannot validly be replaced with +1+1!
@@
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=(-1)^{\pm d}\psi_{n}(e^{\pm i\pi}\eta;e^{\mp i\pi}\textbf{k}_{1},\dots,e^{\mp i\pi}\textbf{k}_{n};\Omega)\,$, (6.11)
@@
$\mathbf{RR}:\psi_{n}^{*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=(-1)^{d}\psi_{n}(\eta;e^{-i\pi}\textbf{k}_{1},\dots,e^{-i\pi}\textbf{k}_{n};e^{i\pi}\Omega)\,.$ (6.12)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=$ $\psi_{n}(e^{-i\pi}\eta;-\textbf{y}_{1},\dots,-\textbf{y}_{n};e^{-i\pi}\Omega)\,;$ (6.13)
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=$ $\psi_{n}(e^{\pm i\pi}\eta;e^{\pm i\pi}\textbf{y}_{1},\dots,e^{\pm i\pi}\textbf{y}_{n};\Omega)\,;$ (6.14)
$\mathbf{RR}:\psi_{n}^{*}(\eta;\textbf{y}_{1},\dots,\textbf{y}_{n};\Omega)=$ $\psi_{n}(\eta;-e^{i\pi}\textbf{y}_{1},\dots,-e^{i\pi}\textbf{y}_{n};e^{-i\pi}\Omega)\,.$ (6.15)
@@
$\mathbf{CRT}:\psi_{n}^{*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=$ $\psi_{n}(e^{-i\pi}\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};e^{-i\pi}\Omega)\,;$ (6.16)
$\mathbf{D}_{-1}^{\pm}:\psi_{n}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=$ $e^{\pm i\pi d}\,\psi_{n}(e^{\pm i\pi}\eta;e^{\mp i\pi}\textbf{k}_{1},\dots,e^{\mp i\pi}\textbf{k}_{n};\Omega)\,;$ (6.17)
$\mathbf{RR}:\psi_{n}^{*}(\eta;\textbf{k}_{1},\dots,\textbf{k}_{n};\Omega)=$ $e^{i\pi d}\,\psi_{n}(\eta;e^{-i\pi}\textbf{k}_{1},\dots,e^{-i\pi}\textbf{k}_{n};e^{-i\pi}\Omega)\,.$ (6.18)